OpenStudy (anonymous):

"If Tom buys a red skateboard, for $50 and then Anne buys the same skateboard for 25% off. How much did Anne pay for her skateboard?

OpenStudy (anonymous):

well so she didn't pay 25% of 50$ and he payed (100% - 25% = 75%) of 50$

OpenStudy (anonymous):

\[\frac{ 75 }{ 100 } \times $50 = \frac{ 75 }{ 2 } \]

OpenStudy (anonymous):

so she only payes 37.5 dollars